# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

### v3

@@ 5.3 Decorated Merge Trees: Function Valued Weights
  • $d_{E}$ is an adaptation of metric here induced to compare merge trees represented by means of editable graphs, essentially truncating the edge at infinity (with some additional details, see Pegoraro (2024a));
  • #$\#T_{f}$ being the cardinality of the set of vertices of $T_{f}$.
@@
Note that Equation 1 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees. The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions. Moreover, the $1$-Wasserstein distance between persistence diagrams satisfies and analogous relationship w.r.t. the bottleneck distances.
As a consequence, our metric is better at picking up global differences and not just local differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams). On the other hand, this also means that is it more sensitive to noise. For more details and examples see Pegoraro (2024a).
[…]
  1. 1.
    diagram or Betti numbers decorated merge trees, on the same line of thought of Curry et al. (2022, 2023b);
  2. 2.
    merge trees decorated with the measure of sublevel sets.
We illustrate this with two qualitative examples: first we consider the datum of a point cloud a then a function defined on a open subset of $\mathbb{R}^{n}$.
$Y_{t}^{k}:=\cup_{p\in Y^{k}}B_{t}(p)$
$B_{t}(p)$ being the metric ball centred in $p$, which is also called the offset filtration. Any point $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$ is a couple $(t,U)$, with $t\in\mathbb{R}$ and $U$ being a path connected component of $Y_{t}^{k}$. Then, one collect some topological information about $U$ and consider a function $f$ with values in $X=\mathbb{R}^{N}_{\geq 0}$, for some $N<n$, defined by:
$f((t,U))=(\beta_{0}(U),\beta_{1}(U),\ldots,\beta_{N}(U)),$
with $\beta_{i}(U)=\dim\text{H}_{i}(U,\mathbb{R})$ being the $i$-th Betti number of $U$ clearly $\beta_{0}(U)=1$. For an example see Pegoraro (2024b). An alternative and richer way to capture shape-related information about $U$ is to consider the point cloud: $U_{k}:=Y^{k}\cap U$ and then build $\text{PD}_{q}(U_{k})$, the persistence diagram of the offset filtration of $U_{k}$ using $q$-dimensional homology. Consequently, one can define $f$ with values in $X=D(\mathbb{R}^{2}_{\Delta},\Delta)$, the space of persistence diagrams with the $1$-Wasserstein distance, defined by:
$f((t,U))=\text{PD}_{q}(U_{k}).$
Consider now a Lebesgue measurable function $g:W\rightarrow\mathbb{R}$, with $W$ open in $\mathbb{R}^{n}$. Build $Y_{t}=f^{-1}((-\infty,t])$, the sublevel set filtration of $f$. And consider $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$. This time another shape-related information that we can estract from a couple $(t,U)\in D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$, is the Lebesgue measure of $U$. And thus define:
$f((t,U))=\mathcal{L}(U),$
with $\mathcal{L}(U)$ being the Lebesgue measure of $U$. For an example of situations that can be tackled with this approach see Figure 4(c) and Figure 4(d). Note that the representation of functional data obtained with the volume-decorated merge trees, lies in between information captured by $L_{p}$ metrics and purely topological information: with merge trees one would not be able to distinguish the two functions in Figure 4(c) and Figure 4(d) (they have the same merge trees), but with $L_{p}$ metrics one would not capture any of the shape similarities between the functions. Moreover, functional data analysis approaches capturing the shape of functions - e.g. aligning functions with re-parametrizations, see for instance the Special Section on Time Warpings and Phase Variation on the Electronic Journal of Statistics, Vol 8 (2) - are difficult to develop over $\mathbb{R}^{n}$, $n\geq 2$.
[…]
$d_{E}(T_{f},T_{g})\leq(K+\max\{\text{diam}Y,\text{diam}Y^{\prime}\})\cdot 4\cdot\#Y\cdot\#Y^{\prime}\cdot d_{G-H}(Y,Y^{\prime}),$
with $d_{H}(Y,Y^{\prime})$ being the Gromov-Hausdorff distance between compact metric spaces. We believe that the first two approaches have the potential to be powerful shape summaries to distinguish between point clouds, as also shown with several examples by Curry et al. (2022, 2023b), employing similar (though not metrically equivalent) tools. Instead, for more details and examples on the use of functional weights for editable graphs, see Pegoraro (2024b). Note that between the metric $d_{E}$, applied to this context, and the metric developed in Curry et al. (2022), hold similar considerations to the ones we made comparing $d_{E}$ and the interleaving distance between merge trees.
(a) A point cloud (left), with the thickening giving the off-set filtration at some particular value (dashed horizontal cyan line) and the merge tree of the filtration (right). At the chosen value the filtration has three path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram.
(c) A real valued function defined on a square with four distinct local minima, associated to downward peaks with different shapes. The associated decorated merge tree (right) accordingly shows four branches (black) and, in red, the (qualitative) volume of the associated downward peaks included in the sublevel sets, which increases with the value of the filtration.
(d) Same situation as in Figure 4(c), but for a different image. We clearly see that the four downward peaks have very different sizes, and that is reflected by the very different volume profiles attached to the branches of the merge tree. Note that the two images have the same (non-decorated) merge trees.
Figure 4: Plots related to Section 5.3.
5.4 Persistent Sets with Embedded Posets: Vector Valued Weights
Here we consider the case of a persistent set $F:P\hookrightarrow\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and focus on the approach described by item 2. in Section 5.2. Where $P$ is a finite poset and the poset structure in $\mathbb{R}^{n}$ is induced by $r=(r_{1},\ldots,r_{n})\leq r^{\prime}=(r^{\prime}_{1},\ldots,r^{\prime}_{n})$ if $r_{i}\leq r^{\prime}_{i}$ for all $i$. This is the kind of situations that works studying multiparameter persistence care about, by making finiteness assumptions like finitely presentable persistence modules (Oudot and Scoccola, 2024) - that is, modules on $\mathbb{R}^{n}$ which are determined by their values on a finite sub-poset of $\mathbb{R}^{n}$. We associate to $F$ a unique editable graph built in this way:
  • we take $P$, build the associated DAG and take its transitive reduction $\mathcal{G}(P)$;
  • build the weight function $\varphi_{\mathcal{G}(P)}$ as in Section 5.2, item 2), so that each edge between $a,b\in\mathbb{R}^{n}_{\geq 0}$, with $a\leq b$ we associate the vector $b-a$;
  • finally we build the display poset of $F_{\mid P}$, which we call $D_{F}$, take its transitive reduction $\mathcal{G}(D_{F})$ and build the weight function $\varphi_{\mathcal{G}(D_{F})}$ by pushing forward $\varphi_{\mathcal{G}(P)}$. See Figure 5 for a detailed example.
We believe that the metric $d_{E}$ is a natural metric to employ to compare these kind of graph representations for multidimensional persistent sets. In fact, the metric exhibits some natural behaviour, for instance as in Figure 6 and, in case $n=1$, and has proven stability results - which we reported in previous sections.
(a) A finite sub-poset of $\mathbb{R}^{2}$.
(b) The transitive reduction of the poset in Figure 5(a).
(c) The DAG associated to a persistent set on the poset in Figure 5(a). Arrows are colored consistently with the ones in the poset in Figure 5(a) and represent the weights associated to the corresponding edges. The maps between sets are represented by colors and follow (row-wise and column-wise) the elder rule.
(d) The editable graph associared to the transitive reduction of the DAG in Figure 5. As in Figure 5, arrows represent weights in $\mathbb{R}^{2}_{\geq 0}$
Figure 5: Plots related to Section 5.4.
(a) Another DAG representing a persistent set over a poset in $\mathbb{R}^{2}$.
(b) Starting from the DAG in Figure 6(a), with two ghosting edits, we can replace the two paths around the square associated to the pink sets, with the two diagonal edge. Thus the DAG in Figure 6(a) and this one are identified by our edit distance.
Figure 6: Additional plots related to Section 5.4.
5.5 Problems to be Faced
In the previous sections we have laid out some ideas which suggest how different persistent sets can be represented via editable graphs with weight functions taking values in different editable spaces. As pointed out at the beginning of Section 5, these ideas present some shortcomings which need to be overcome and some formal results which need to be established in order for the ideas to be properly applied. In this section we present the major issues affecting the proposed approach: some of them are already tackled and solved in other works, while other are left open and serve as invitations for future investigations.
In the same way a point cloud is represented by the pairwise distances between its points only up to isometry, similarly, collecting distances or vectors as weights for the edges does not provide a map from persistent sets to graphs that can be inverted up to natural isomorphisms. Consider the case of a one dimensional persistent set $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$ and define for every $h\in\mathbb{R}$ the filtration $Y_{h}(t):=Y(t+h)$. If we then proceed by building the display poset and obtaining the associated graph as discussed in Section 5.4, then the editable graphs associated to $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are going to be isomorphic as editable graphs, as their posets are going to be isomorphic as posets and $r\mapsto r+h$ is an isometry. While the persistent sets $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are not isomorphic as functors. And this is true also for $Y:\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and $h\in\mathbb{R}^{n}$.
As already mentioned, the fiber of the map sending a constructible persistent set $S$ into the editable graph $\mathcal{G}(D_{S_{\mid C}})$ (up to isomorphism) contains a lot of persistent sets which are not isomorphic. Moreover, we believe that, for any representation of persistent set/persistence module, it is always advisable to produce some stability result, to ensure the interpretability of the pipeline.
In the particular case of $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$, both the issues presented have been taken care in Pegoraro (2024a). For all other scenarios these problems still await formal investigations.
Additionally, we would like also to extend this framework to Reeb graphs, similarly to what was done for merge trees. This is motivated by the extent to which Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018) and used in applications (Nielson et al., 2015; Rizvi et al., 2017). They add another layer of difficulties as a metric being insensitive to topological changes as the ones depicted in Figure 6 it is not desirable. This could be solved by modeling the possibility of having multiple edges connecting two vertices. It would also be important also to develop an algorithm to compute or approximate $d_{E}$ which works for Reeb graphs.
6 Trees
In the remaining part of the manuscript we specialize our discussion on a particular kind of directed graphs: trees. Thanks to 6 we have seen that this scenario still includes a lot of interesting situations, covering all constructible (one dimensional) persistent sets whose filtration which definitively is path-connected. Moreover, the peculiar structure of trees allows us to prove some key results which greatly decrease the complexity of computing optimal mappings between two trees, resulting in a bottom-up algorithm which is implemented and tested. We leave to future works the investigation of algorithms for more complicated kinds of graphs.
Remark 5
From now on our point of view is predominantly computational, thus we always assume to have finite graphs.
Definition 18
A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$. We indicate the root of the tree with $r_{T}$. The order of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the extremes, and is called $ord_{T}(v)$. Any vertex with an edge connecting it to the root is its child and the root is its father: this is the first step of a recursion which defines the father and children relationship for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$. The relation $child<father$ generates a partial order on $V_{T}$ which induces an orientation on the edges. The edges in $E_{T}$ are identified in the form of ordered couples $(a,b)$ meaning $a<b$ and $a\rightarrow b$. A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to an editable tree by calling it dendrogram, as formalized by the following definition.
Definition 19
A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets. Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation. Moreover, since we now have a canonical way to associate edges to vertices, we can state the definition of mappings in a more directed and handy way, associating an edit operation to every edge/vertex in a tree - a part from the root. Such association is not available in general for graphs. In this way the ghosting of an edge can be described via the removal of individual vertices, and similarly deletions and shrinkings can be referred to as applied on vertices instead of edges. With an abuse of notation we call mappings the following objects, which from now on replace the mappings defined in Section 4.2.
The letters "$D"$ and "$G"$ are used to indicate deletion and ghosting. A mapping between $T$ and $T^{\prime}$ is thus a set $M\subset(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})$ satisfying:
  • (M1)
    consider the projection of the Cartesian product $(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})\rightarrow(E_{T}\cup\{"D","G"\})$; we can restrict this map to $M$ obtaining $\pi_{T}:M\rightarrow(E_{T}\cup\{"D","G"\})$. The maps $\pi_{T}$ and $\pi_{T^{\prime}}$ are surjective on $E_{T}\subset(E_{T}\cup\{"D","G"\})$ and $E_{T^{\prime}}\subset(E_{T^{\prime}}\cup\{"D","G"\})$ respectively;
  • (M2)
    $\pi_{T}$ and $\pi_{T^{\prime}}$ are injective;
  • (M3)
    $M\cap(V_{T}\times V_{T^{\prime}})$ is such that, given $(a,b)$ and $(c,d)\in M\cap(V_{T}\times V_{T^{\prime}})$, $a>c$, if and only if $b>d$;
  • (M4)
    if $(a,"G")\in M$ (or analogously ("$G",a)$), then after applying all deletions of the form $(v,"D")\in M$, the vertex $a$ becomes an order $2$ vertex. In other words: let $child(a)=\{b_{1},..,b_{n}\}$. Then there is one and only one $i$ such that for all $j\neq i$, for all $v\in V_{sub(b_{j})}$, we have $(v,"D")\in M$; and there is one and only one $c$ such that $c=\max\{x<b_{i}\mid(x,y)\in M$ for any $y\in V_{T^{\prime}}\}$.
We highlight some points with the following remarks.
Remark 6
This definition is equivalent to the one given in Section 4.2. In fact: (M1) and (M2) are equivalent to (G1) plus the fact that vertices not appearing in the image of $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{T}$ are deleted. Property (G2) is equivalent to (M4): if a segment $s$ which is not simply an edge, and so it induces some ghostings - is such that all the edges in its open star are either deleted or belong to $Im(s)$, this means that all the vertices in $s^{o}$ are of order $2$ after the deletions. Lastly, (M3) implies that the trees considered are isomorphic as directed graphs once all deletions and ghostings are applied on the starting and target trees.
Remark 7
In literature usually mappings are subsets of $E_{T}\times E_{T^{\prime}}$ and their properties are equivalent to (M2) and (M3) (see Hong et al. (2017); Sridharamurthy et al. (2020); Wetzels et al. (2022) and references therein). In this way, mappings coherently maps edges into edges, respecting the tree ordering, inducing in this way the needed relabeling operations - see also 3. The differences with our definition are caused by the introduction of the ghosting edits. Accordingly, the way in which mappings parametrize edit paths in the present work is completely novel and, for instance, it establishes some (partial) ordering between the edits. Note that the set of mappings we consider is bigger than in usual edit distances - thus computing the distance has a higher computational cost.
6.1 Computing the Edit Distance for Trees: Decomposition Properties
In these last sections we develop an algorithm to compute the edit distance for dendrograms.
@@ 6.1 Computing the Edit Distance for Trees: Decomposition Properties
Name $T_{2}$ the only representative without order $2$ vertices inside the equivalence class of $T$. One can always suppose that a dendrogram is given without order $2$ vertices. Thus, for notational convenience, from now on we suppose $T=T_{2}$ and $T^{\prime}=T^{\prime}_{2}$. To help us in the calculations define a particular subset of mappings $M_{2}(T,T^{\prime})\subset\operatorname*{Mapp}(T,T^{\prime})$.
Definition 20
A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has maximal ghostings if $(v,"G")\in M$ if and only if $v$ is of order $2$ after the deletions in $T$ and, similarly ("$G",w)\in M$ if and only if $w$ is of order $2$ after the deletions in $T^{\prime}$. A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has minimal deletions if $(v,"D")\in M$ only if neither $v$ nor $father(v)$ are of order $2$ after applying all the other deletions in $T$ and, similarly, ("$D",w)\in M$ only if $w$ is not of order $2$ after applying all the other deletions in $T^{\prime}$.
We collect all mappings with maximal ghostings in the set $M_{\mathcal{G}}(T,T^{\prime})$ and paths with minimal deletions in the set $M_{\mathcal{D}}(T,T^{\prime})$. Lastly: $M_{2}(T,T^{\prime}):=M_{\mathcal{G}}(T,T^{\prime})\bigcap M_{\mathcal{D}}(T,T^{\prime})$.
In other words with mappings in $M_{2}(T,T^{\prime})$ we are always eliminating all the order $2$ vertices which arise from deletions and we are not deleting edges which we can shrink. The following lemma then applies.
Lemma 1
$\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}=\min\{cost(M)\mid M\in M_{2}(T,T^{\prime})\}$
In addition to that, we consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows. Recall that, using $E_{T}\cong V_{T}-\{r_{T}\}$, we can induce $\pi_{T}:E_{T}\times E_{T^{\prime}}\rightarrow V_{T}$.
@@
    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $father>son$. This means that any two distinct vertices of $T$ (respectively of $T^{\prime}$) which appear in $M^{*}$ are incomparable with respect to “$>$”;
@@
Consider now $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. Starting from such set of couples we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each couple $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices. The vertices below $x$ and $y$ will be taken care separately. See also Figure 7 for a visual example.
@@
And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$. For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$. Note that $\zeta_{x}^{\Omega(x)}$ is a segment (see 9).
@@
    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,"D")\in\alpha(M^{*})$;
@@
    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,"D")\in\alpha(M^{*})$;
@@
      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,"G")\in\alpha(M^{*})$
@@
      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,"D")\in\alpha(M^{*})$;
@@
The idea behind $\alpha(M^{*})$ is that, for all couples $(x,y)\in M^{*}$, we want to make the ghostings to turn the segments $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other. As we already anticipated, $\alpha(M^{*})$ takes care of all the vertices in $V_{T}$ and $V_{T^{\prime}}$, a part from the sets $\cup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\cup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. For this reason we say that $\alpha(M^{*})$ is a partial mapping.
@@
Figure 7: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $G_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T_{y})$, as in the algorithm in Section 6.2. The set $M^{*}=\{(v,w),(v^{\prime},w^{\prime}),(v^{\prime\prime},w^{\prime\prime})\}$ satisfies (A1),(A2). The set $\alpha(M^{*})$ is made by the deletions/insertions indicated by the red edges, the ghostings/splittings indicated by the yellow and the shrinkings given by edges of the same color, different from red. To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 7 all the black vertices covered by shaded regions must be deleted.
6.2 Computing the Edit Distance for Trees: Dynamical Binary Linear Programming problems

### v4

@@ 5.3 Stability Properties
  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between merge trees;
  • $\text{size}(T_{f})$ being the cardinality of the edge set of $T_{f}$.
@@
Equation 2 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees. The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions. Moreover, we have already pointed out that the $1$-Wasserstein distance between persistence diagrams satisfies an analogous relation w.r.t. the bottleneck distances:
$d_{B}(D_{f},D_{g})\leq W_{1}(D_{f},D_{g})\leq(\#D_{f}+\#D_{g})\parallel f-g\parallel_{\infty}.$
As a consequence, our metric is better at picking up differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams). On the other hand, this also means that is it more sensitive to noise. Building on these considerations, in Pegoraro (2024a) it is argued that such stability properties may be preferable to universal stability properties, in which the operator from functions to topological representation is $1$-Lipschitz. Moreover, such comparison is framed in analogy to the bias-variance tradeoff in statistical modeling. We further observe that, since merge trees can be viewed as a special case of Reeb graphs, the same considerations we made for $d_{I}$ also apply to all the metrics discussed in Bollen (2022)—including those introduced in Bauer et al. (2014); Di Fabio and Landi (2016); Bauer et al. (2020)—due to their metric equivalence with the interleaving distance.
Lastly, Pegoraro (2024b) considers dendrograms enriched with Betti numbers of (path-connected components of) sublevel sets of functions, similarly to Equation 1. It is therein shown that:
$d_{E}(\varphi_{f}^{\Theta_{p}},\varphi_{g}^{\Theta_{p}})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))2C\parallel f-g\parallel_{\infty},$
with:
  • $\varphi_{f}^{\Theta_{p}}$ and $\varphi_{g}^{\Theta_{p}}$ are the dendrograms associated to (enriched) the persistent sets of the sublevel set filtrations of $f$ and $g$;
  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between functions defined on display posets of persistent sets;
  • $C>0$ is an upper bound to the values of the Betti numbers;
  • $\text{Crit}(f)$ is the set of values where the homology of the sublevel sets of $f$ changes, in any dimension.
To conclude, we note that, if we consider the persistent homology of the sublevel set filtration of $f$ and $g$, in dimensions $i=0,\ldots,N$, and employ the persistence diagrams $D_{f}^{i}$ and $D_{g}^{i}$, we have:
$\sum_{i=0}^{N}W_{1}(D^{i}_{f},D^{i}_{g})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))\parallel f-g\parallel_{\infty}.$
6 Computing the Edit Distance
In these last section we develop an algorithm to compute the edit distance for dendrograms.
6.1 Decomposition Properties
@@
Name $T_{2}$ the only representative without degree $2$ vertices inside the equivalence class of $T$. One can always suppose that a dendrogram is given without degree $2$ vertices. Thus, for notational convenience, from now on we suppose $T=T_{2}$ and $T^{\prime}=T^{\prime}_{2}$.
We consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows. Recall that, using $E_{T}\cong V_{T}-\{r_{T}\}$, we can induce $\pi_{T}:E_{T}\times E_{T^{\prime}}\rightarrow V_{T}$.
@@
    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $parent>child$. This means that any two distinct vertices of $T$ (respectively of $T^{\prime}$) which appear in $M^{*}$ are incomparable with respect to “$>$”;
@@
Consider now $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. Starting from such set of pairs we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each pair $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices. The vertices below $x$ and $y$ will be taken care separately. See also Figure 5 for a visual example.
@@
And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$. For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$.
@@
    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,\mathfrak{D})\in\alpha(M^{*})$;
@@
    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
@@
      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,\mathfrak{G})\in\alpha(M^{*})$
@@
      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
@@
The idea behind $\alpha(M^{*})$ is that, for all pairs $(x,y)\in M^{*}$, we want to make the ghostings to turn the paths $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other. As we already anticipated, $\alpha(M^{*})$ takes care of all the vertices in $V_{T}$ and $V_{T^{\prime}}$, a part from the sets $\cup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\cup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. For this reason we say that $\alpha(M^{*})$ is a partial mapping.
@@
Figure 5: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $T^{\prime}_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T^{\prime}_{y})$, as in the algorithm in Section 6.2. The set $M^{*}=\{(v,w),(v^{\prime},w^{\prime}),(v^{\prime\prime},w^{\prime\prime})\}$ satisfies (A1),(A2). The set $\alpha(M^{*})$ is made by the deletions/insertions indicated by the red edges, the ghostings/splittings indicated by the yellow and the shrinkings given by edges of the same color, different from red. To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 6 all the black vertices covered by shaded regions must be deleted.
6.2 Dynamical Binary Linear Programming problems